### v7

@@ 4.1 Preliminary Lemmas
$\mathbb{E}\lVert{y_{n}(t)}\rVert_{L^{q}}^{p}\leqslant\lVert{v_{n}}\rVert_{L^{q}}^{p}+c\,\mathbb{E}\bigg{[}\int_{0}^{t}1+\lVert{y_{n}(s)}\rVert_{L^{q}}^{p}\,\mathrm{d}s\bigg{]},\quad\forall t\in[0,T].$
Applying Gronwall’s lemma to the above inequality yields
$\lVert{y_{n}}\rVert_{C([0,T];L^{p}(\Omega;L^{q}))}\leqslant c\big{(}1+\lVert{v_{n}}\rVert_{L^{q}}\big{)},$
which implies the desired inequality Eq. 4.7. This completes the proof. ∎
Proof of Theorem 3.1. We split the proof into the following two steps.
Step 1. We proceed to show that equation Eq. 1.1 has a unique mild solution $y$, and that Eq. 3.1 is satisfied. Let $p,q\in(2,\infty)$ be fixed but arbitrary. For each $n\in\mathbb{N}$, denote by $y_{n}$ the mild solution of equation Eq. 4.6, where $v_{n}$ is defined as
$v_{n}:=\mathop{\mathrm{argmin}}_{w\in\mathrm{span}\{\phi_{0},\ldots,\phi_{n}\}}\lVert{v-w}\rVert_{L^{3q}}.$
For any $m,n\in\mathbb{N}$, put $e_{m,n}:=y_{m}-y_{n}$. By definition, we infer that
$\mathrm{d}e_{m,n}(t)=(\Delta e_{m,n}+e_{m,n}+y_{n}^{3}-y_{m}^{3})(t)\,\mathrm{d}t+\big{(}F(y_{m}(t))-F(y_{n}(t))\big{)}\,\mathrm{d}W_{H}(t),\quad t\in[0,T].$
Observing that the integral over $\mathcal{O}$ of $\lvert{e_{m,n}(t)}\rvert^{q-2}e_{m,n}(t)(y_{n}^{3}-y_{m}^{3})(t)$ is non-positive for all $t\in[0,T]$, and following the proof of Eq. 4.7, we can deduce that
$\lVert{e_{m,n}}\rVert_{L_{\mathbb{F}}^{3p}(\Omega\times(0,T);L^{3q})}\leqslant c\lVert{v_{m}-v_{n}}\rVert_{L^{3q}}.$ (4.14)
On the other hand, from the fact that $\dot{H}^{2,2}$ is dense in $L^{3q}$ and that $(\phi_{n})_{n\in\mathbb{N}}$ is a complete orthogonal system in $\dot{H}^{2,2}$, we conclude that $(v_{n})_{n\in\mathbb{N}}$ is a Cauchy sequence in $L^{3q}$, converging to $v$. Consequently, there exists a unique process $y\in L_{\mathbb{F}}^{3p}(\Omega\times(0,T);L^{3q})$ such that
$\lim_{n\to\infty}y_{n}=y\quad\text{ in }L_{\mathbb{F}}^{3p}(\Omega\times(0,T);L^{3q}).$ (4.15)
@@
$\lVert{y_{n}^{3}-y^{3}}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}$
@@
$\lim_{n\to\infty}y_{n}^{3}=y^{3}\quad\text{ in }L_{\mathbb{F}}^{p}(\Omega\times(0,T);L^{q}),$
@@
$\lim_{n\to\infty}y_{n}-y_{n}^{3}=y-y^{3}\quad\text{ in }L_{\mathbb{F}}^{p}(\Omega\times(0,T);L^{q}).$
Using Lemma 4.3, this directly implies the following convergence of $S_{0}(y_{n}-y_{n}^{3})$:
$\lim_{n\to\infty}S_{0}(y_{n}-y_{n}^{3})=S_{0}(y-y^{3})\quad\text{ in }L_{\mathbb{F}}^{p}(\Omega\times(0,T);L^{q}).$
Furthermore, invoking Eqs. 4.4 and 4.15 yields
$\lim_{n\to\infty}F(y_{n})=F(y)\quad\text{ in }L_{\mathbb{F}}^{p}(\Omega\times(0,T);\gamma(H,L^{q})),$
and subsequently by Lemma 4.4 we infer the following convergence for $S_{1}F(y_{n})$:
$\lim_{n\to\infty}S_{1}F(y_{n})=S_{1}F(y)\quad\text{ in }L_{\mathbb{F}}^{p}(\Omega\times(0,T);L^{q}).$
@@
Additionally, applying Lemma 4.2 with $\alpha=\beta=0$ and noting the convergence $\lim_{n\in\infty}\lVert{v_{n}-v}\rVert_{L^{q}}=0$, we establish the following convergence for $G_{n}$:
$\lim_{n\to\infty}G_{n}=G\quad\text{ in }L^{p}(0,T;L^{q}),$
@@
where $G$ is defined by Eq. 2.10, and $G_{n}(t):=e^{t\Delta}v_{n}$ for each $t\in[0,T]$. Finally, combining the derived convergences of $y_{n}$, $G_{n}$, $S_{0}(y_{n}-y_{n}^{3})$, and $S_{1}F(y_{n})$, and recalling that $y_{n}=G_{n}+S_{0}(y_{n}-y_{n}^{3})+S_{1}F(y_{n})$, we obtain by passing to the limit $n\to\infty$ that
$y=G+S_{0}(y-y^{3})+S_{1}F(y)\quad\text{in }L_{\mathbb{F}}^{p}(\Omega\times(0,T);L^{q}).$
This confirms that $y$ is a mild solution of Eq. 1.1, and it is evident that Eq. 3.1 holds. To prove the uniqueness of the mild solution, assume $\widetilde{y}$ is another mild solution. Consider the difference $e=y-\widetilde{y}$. Then $e$ satisfies the equation
@@
$\begin{cases}\mathrm{d}e(t)=(\Delta e+e+\widetilde{y}^{3}-y^{3})(t)\,\mathrm{d}t+\big{(}F(y(t))-F(\widetilde{y}(t))\big{)}\,\mathrm{d}W_{H}(t),&t\in[0,T],\\ e(0)=0.\end{cases}$
By an argument similar to that used in Eq. 4.14, we find that
$\lVert{e}\rVert_{L^{2}(\Omega\times(0,T);L^{2})}\leqslant 0,$
which implies that $e=0$ almost surely, confirming the uniqueness of the mild solution to Eq. 1.1.
@@
Step 2. Fix any $p\in(2,\infty)$ and $q\in[2,\infty)$. By virtue of the regularity assertion in Eq. 3.1 and the growth estimate Eq. 4.1 for $F$, we deduce that $y-y^{3}$ belongs to $L_{\mathbb{F}}^{p}(\Omega\times(0,T);L^{q})$ and $F(y)$ is an element of $L_{\mathbb{F}}^{p}(\Omega\times(0,T);\gamma(H,L^{q}))$. Hence, the application of Lemma 4.3, in conjunction with Lemma 4.1, gives that
@@
while Lemma 4.4 implies $S_{1}F(y)\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{1,q})$. Additionally, Lemma 4.2 ensures that $G\in L^{p}(0,T;\dot{H}^{2/p-\epsilon,q})$ for all $\epsilon>0$. Consequently, by the mild solution representation Eq. 2.9, we arrive at
@@
Finally, combining Eqs. 2.9, 4.16 and 4.17 establishes the regularity assertions in Eqs. 3.4 and 3.5. Moreover, the regularity assertion in Eq. 3.2 follows directly from Eq. 3.4 and the observation that $G\in C([0,T];L^{q})$. This completes the proof of Theorem 3.1.
@@
We begin by summarizing several well-established results. As known from [8], it holds that
$\|P_{h}\|_{\mathcal{L}(L^{q})}\text{ is uniformly bounded with respect to }h\text{ for all }q\in(1,\infty).$ (4.18)
Consequently, the operator $P_{h}$ possesses the following approximation property:
@@
Inequality (2.5) from [1], combined with the observation that
$\lVert{\Delta_{h}^{-1}P_{h}\Delta}\rVert_{\mathcal{L}(\dot{H}^{1,2})}\leqslant 1,$
yields, through interpolation, that for any $q\in[2,\infty)$, the operator norm
$\lVert{\Delta_{h}^{-1}P_{h}\Delta}\rVert_{\mathcal{L}(\dot{H}^{1,q})}$
is uniformly bounded with respect to $h$. Utilizing this fundamental stability estimate, the approximation property Eq. 4.19, the interpolation theory as stated in [22, Theorem 2.6], and the fact that $\Delta^{-1}$ is a bounded linear operator mapping from $\dot{H}^{-s,q}$ to $\dot{H}^{2-s,q}$ for all $s\in[0,2]$, standard arguments then lead to the conclusion that
$\|P_{h}\Delta^{-1}-\Delta_{h}^{-1}P_{h}\|_{\mathcal{L}(\dot{H}^{\beta-2,q},\dot{H}^{\alpha,q})}\leqslant ch^{\beta-\alpha}\quad\text{for all }0\leqslant\alpha\leqslant 1\leqslant\beta\leqslant 2.$ (4.20)
Let $e^{t\Delta_{h}}$, for $t\in[0,\infty)$, denote the analytic semigroup generated by the discrete Laplace operator $\Delta_{h}$. Using the established estimates for the resolvents of $\Delta_{h}$ from [1] and applying (2.128) from [38], we obtain the well-known growth estimate: for any $\alpha\in[0,1]$, $q\in(1,\infty)$, and $t\in(0,T]$,
@@
Furthermore, utilizing techniques analogous to those found in Theorem 6.10 and Lemma 20.7 of [35], one can readily derive the following well-known inequality: for any $\alpha\in[0,2]$, $q\in(1,\infty)$, and $t\in[0,T]$,
@@
For the special case where $q=2$, the reader is referred to Theorem 3.2 of [35] for a detailed treatment of this result.
Let $q\in(1,\infty)$. For any $\alpha\in\mathbb{R}$, denote by $\dot{H}_{h}^{\alpha,q}$ the space $X_{h}$ equipped with the norm
$\|v_{h}\|_{\dot{H}_{h}^{\alpha,q}}:=\|(-\Delta_{h})^{\alpha/2}v_{h}\|_{L^{q}}\quad\text{for all }v_{h}\in X_{h}.$
According to Theorem 3.1 (ii) in [17], the boundedness of the purely imaginary powers of $-\Delta_{h}$ in $\mathcal{L}(\dot{H}_{h}^{0,q})$ is independent of the spatial mesh size $h$. Consequently, by applying Theorem 4.17 from [22], we deduce that for any $\alpha\in(0,2)$,
$c_{0}\|v_{h}\|_{\dot{H}_{h}^{\alpha,q}}\leqslant\|v_{h}\|_{[\dot{H}_{h}^{0,q},\dot{H}_{h}^{2,q}]_{\alpha/2}}\leqslant c_{1}\|v_{h}\|_{\dot{H}_{h}^{\alpha,q}}\quad\text{for all }v_{h}\in X_{h},$ (4.23)
where $c_{0}$ and $c_{1}$ are positive constants independent of $h$. Here, $[\dot{H}_{h}^{0,q},\dot{H}_{h}^{2,q}]_{\alpha/2}$ denotes the complex interpolation space between $\dot{H}_{h}^{0,q}$ and $\dot{H}_{h}^{2,q}$ (see [22, Chapter 2]). This result is important as it enables the application of interpolation theory to the space $\dot{H}_{h}^{s,q}$ for all $s\in\mathbb{R}$. For example, given the standard inverse estimate
$\|v_{h}\|_{\dot{H}_{h}^{2,q}}\leqslant ch^{-2}\|v_{h}\|_{L^{q}}\quad\text{for all }v_{h}\in X_{h},$
we can invoke [22, Theorem 2.6] to derive the following inverse estimate: for any $0\leqslant\alpha<\beta<\infty$,
$\|v_{h}\|_{\dot{H}_{h}^{\beta,q}}\leqslant ch^{\alpha-\beta}\|v_{h}\|_{\dot{H}_{h}^{\alpha,q}}\quad\text{for all }v_{h}\in X_{h}.$
Furthermore, for any $\alpha\in[0,1]$, there exist positive constants $c_{0}$ and $c_{1}$, independent of $h$, such that
$c_{0}\|v_{h}\|_{\dot{H}^{\alpha,q}}\leqslant\|v_{h}\|_{\dot{H}_{h}^{\alpha,q}}\leqslant c_{1}\|v_{h}\|_{\dot{H}^{\alpha,q}}\quad\text{for all }v_{h}\in X_{h}.$
Henceforth, the aforementioned theoretical results will be utilized without further explicit reference.
We now introduce a discrete deterministic maximal $L^{p}$-regularity estimate and its stochastic counterpart. For a function $g_{h}\in L^{1}(0,T;\dot{H}_{h}^{0,2})$, the discrete deterministic convolution operator $S_{0,h}$ is defined as
$(S_{0,h}g_{h})(t):=\int_{0}^{t}e^{(t-s)\Delta_{h}}g_{h}(s)\,\mathrm{d}s,\quad\text{for }t\in[0,T].$ (4.24)
For a process $g_{h}\in L_{\mathbb{F}}^{2}(\Omega\times(0,T);\gamma(H,\dot{H}_{h}^{0,2}))$, the discrete stochastic convolution operator $S_{1,h}$ is given by
$(S_{1,h}g_{h})(t):=\int_{0}^{t}e^{(t-s)\Delta_{h}}g_{h}(s)\,\mathrm{d}W_{H}(s),\quad\text{for }t\in[0,T].$ (4.25)
The proofs of the following two lemmas can be found in [16] for the first and in [17] for the second.

### v8

@@ 4.1 Preliminary Lemmas
$\mathbb{E}\lVert{y^{(n)}(t)}\rVert_{L^{q}}^{p}\leqslant\lVert{y_{0}^{(n)}}\rVert_{L^{q}}^{p}+c\,\mathbb{E}\bigg{[}\int_{0}^{t}1+\lVert{y^{(n)}(s)}\rVert_{L^{q}}^{p}\,\mathrm{d}s\bigg{]},\quad\forall t\in[0,T].$
An application of Gronwall’s inequality then yields
$\lVert{y^{(n)}}\rVert_{C([0,T];L^{p}(\Omega;L^{q}))}\leqslant c\big{(}1+\lVert{y_{0}^{(n)}}\rVert_{L^{q}}\big{)},$
which implies the desired estimate Eq. 4.7. This completes the proof. ∎
4.2 Proof of Theorem 3.1
We split the proof into the following three steps.
Step 1. We proceed to show that equation Eq. 1.1 has a unique mild solution $y$, and that Eq. 3.1 is satisfied. Let $p\in(2,\infty)$ and $q\in[4,\infty)$ be arbitrarily chosen. For each $n\in\mathbb{N}$, denote by $y^{(n)}$ the mild solution of equation Eq. 4.6, where $y_{0}^{(n)}$ is defined as
$y_{0}^{(n)}:=\mathop{\mathrm{argmin}}_{w\in\mathrm{span}\{\phi_{0},\ldots,\phi_{n}\}}\lVert{y_{0}-w}\rVert_{L^{3q}}.$
For any $m,n\in\mathbb{N}$, put $e_{m,n}:=y^{(m)}-y^{(n)}$. By definition, we infer that
$\mathrm{d}e_{m,n}(t)=(\Delta e_{m,n}+e_{m,n}+(y^{(n)})^{3}-(y^{(m)})^{3})(t)\,\mathrm{d}t+\big{(}F(y^{(m)}(t))-F(y^{(n)}(t))\big{)}\,\mathrm{d}W_{H}(t),\quad t\in[0,T].$
Observing that the integral over $\mathcal{O}$ of
$\lvert{e_{m,n}(t)}\rvert^{3q-2}e_{m,n}(t)[\Delta e_{m,n}(t)+(y^{(n)}(t))^{3}-(y^{(m)}(t))^{3}]$
is non-positive for all $t\in[0,T]$, and following the proof of Eq. 4.7, we can deduce that
$\lVert{e_{m,n}}\rVert_{L_{\mathbb{F}}^{3p}(\Omega\times(0,T);L^{3q})}\leqslant c\lVert{y_{0}^{(m)}-y_{0}^{(n)}}\rVert_{L^{3q}}.$ (4.14)
On the other hand, from the fact that $\dot{H}^{2,2}$ is dense in $L^{3q}$ and that $(\phi_{n})_{n\in\mathbb{N}}$ is a complete orthogonal system in $\dot{H}^{2,2}$, we conclude that $(y_{0}^{(n)})_{n\in\mathbb{N}}$ is a Cauchy sequence in $L^{3q}$ and converges to $y_{0}$ in $L^{3q}$, i.e.,
$\lim_{n\to\infty}\lVert{y_{0}^{(n)}-y_{0}}\rVert_{L^{3q}}=0.$ (4.15)
Consequently, there exists a unique process $y\in L_{\mathbb{F}}^{3p}(\Omega\times(0,T);L^{3q})$ such that
$\lim_{n\to\infty}y^{(n)}=y\quad\text{ in }L_{\mathbb{F}}^{3p}(\Omega\times(0,T);L^{3q}).$ (4.16)
@@
$\lVert{(y^{(n)})^{3}-y^{3}}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}$
@@
$\lim_{n\to\infty}(y^{(n)})^{3}=y^{3}\quad\text{ in }L_{\mathbb{F}}^{p}(\Omega\times(0,T);L^{q}),$
@@
$\lim_{n\to\infty}y^{(n)}-(y^{(n)})^{3}=y-y^{3}\quad\text{ in }L_{\mathbb{F}}^{p}(\Omega\times(0,T);L^{q}).$
Using Lemma 4.3, this directly implies the following convergence of $S_{0}(y^{(n)}-(y^{(n)})^{3})$:
$\lim_{n\to\infty}S_{0}(y^{(n)}-(y^{(n)})^{3})=S_{0}(y-y^{3})\quad\text{ in }L_{\mathbb{F}}^{p}(\Omega\times(0,T);L^{q})\cap L_{\mathbb{F}}^{p}(\Omega;C([0,T];L^{q})).$
Furthermore, invoking Eqs. 4.4 and 4.16 yields
$\lim_{n\to\infty}F(y^{(n)})=F(y)\quad\text{ in }L_{\mathbb{F}}^{p}(\Omega\times(0,T);\gamma(H,L^{q})),$
and subsequently by Lemma 4.4 we infer the following convergence for $S_{1}F(y^{(n)})$:
$\lim_{n\to\infty}S_{1}F(y^{(n)})=S_{1}F(y)\quad\text{ in }L_{\mathbb{F}}^{p}(\Omega\times(0,T);L^{q})\cap L_{\mathbb{F}}^{p}(\Omega;C([0,T];L^{q})).$
@@
Additionally, applying Lemma 4.2 with $\alpha=\beta=0$ and using Eq. 4.15, we establish the following convergence for $G^{(n)}$:
$\lim_{n\to\infty}G^{(n)}=G\quad\text{ in }C([0,T];L^{q}),$
@@
where $G$ is defined by Eq. 2.8, and $G^{(n)}(t):=e^{t\Delta}y_{0}^{(n)}$ for each $t\in[0,T]$. Finally, combining the derived convergences of $y^{(n)}$, $G^{(n)}$, $S_{0}(y^{(n)}-(y^{(n)})^{3})$, and $S_{1}F(y^{(n)})$, and recalling that $y^{(n)}=G^{(n)}+S_{0}(y^{(n)}-(y^{(n)})^{3})+S_{1}F(y^{(n)})$, we obtain by passing to the limit $n\to\infty$ that
$y=G+S_{0}(y-y^{3})+S_{1}F(y)\quad\text{in }L_{\mathbb{F}}^{p}(\Omega\times(0,T);L^{q})\cap L_{\mathbb{F}}^{p}(\Omega;C([0,T];L^{q})).$
This confirms that $y$ is a mild solution of Eq. 1.1, and it is evident that Eq. 3.1 holds.
Step 2. We now consider the uniqueness of the mild solution $y$, which can be proved by a standard localization argument. Suppose $\widetilde{y}\in L_{\mathbb{F}}^{0}(\Omega;C([0,T];L^{4}))$ is another mild solution of Eq. 1.1, and define the difference $e:=y-\widetilde{y}$. Then $e$ satisfies the stochastic evolution equation
@@
$\begin{cases}\mathrm{d}e(t)=\left(\Delta e(t)+e(t)+\widetilde{y}^{3}(t)-y^{3}(t)\right)\mathrm{d}t\\ \qquad\qquad{}+\left(F(y(t))-F(\widetilde{y}(t))\right)\mathrm{d}W_{H}(t)\quad\text{in $\dot{H}^{-1,2}$,}\quad t\in[0,T],\\ e(0)=0.\end{cases}$
For $m\geqslant 1$, define the stopping time
$\sigma_{m}:=\inf\left\{t\in[0,T]\colon\max\left(\|y(t)\|_{L^{4}},\|\widetilde{y}(t)\|_{L^{4}}\right)\geqslant m\right\},$
with $\sigma_{m}=T$ if the set is empty. Set $e_{m}(t):=e(t\wedge\sigma_{m})$ for $t\in[0,T]$. Then $e_{m}$ satisfies the stopped equation
$\begin{cases}\mathrm{d}e_{m}(t)=\mathbbm{1}_{[0,\sigma_{m}]}(t)\left(\Delta e_{m}(t)+e_{m}(t)+\widetilde{y}^{3}(t\wedge\sigma_{m})-y^{3}(t\wedge\sigma_{m})\right)\mathrm{d}t\\ \qquad\qquad{}+\mathbbm{1}_{[0,\sigma_{m}]}(t)\left(F(y(t))-F(\widetilde{y}(t))\right)\mathrm{d}W_{H}(t)\quad\text{in $\dot{H}^{-1,2}$},\quad t\in[0,T],\\ e_{m}(0)=0.\end{cases}$
Applying Itô’s formula to $\|e_{m}(t)\|_{L^{2}}^{2}$ and taking expectation yields, for all $t\in[0,T]$,
$\mathbb{E}\|e_{m}(t)\|_{L^{2}}^{2}$ $=2\mathbb{E}\bigg{[}\mathbbm{1}_{[0,\sigma_{m}]}(s)\big{\langle}\Delta e_{m}(s)+e_{m}(s)+\widetilde{y}^{3}(s)-y^{3}(s),\,e_{m}(s)\big{\rangle}\,\mathrm{d}s\bigg{]}$
$\quad+\mathbb{E}\bigg{[}\int_{0}^{t}\mathbbm{1}_{[0,\sigma_{m}]}(s)\|F(y(s))-F(\widetilde{y}(s))\|_{\gamma(H,L^{2})}^{2}\,\mathrm{d}s\bigg{]},$
[…]
$\mathbbm{1}_{[0,\sigma_{m}]}(s)\big{\langle}\Delta e_{m}(s)+e_{m}(s)+\widetilde{y}^{3}(s)-y^{3}(s),\,e_{m}(s)\big{\rangle}\leqslant\mathbbm{1}_{[0,\sigma_{m}]}(s)\|e_{m}(s)\|_{L^{2}}^{2},$
it follows that
$\mathbb{E}\|e_{m}(t)\|_{L^{2}}^{2}\leqslant\mathbb{E}\bigg{[}\int_{0}^{t}\mathbbm{1}_{[0,\sigma_{m}]}(s)\left(\|e_{m}(s)\|_{L^{2}}^{2}+\|F(y(s))-F(\widetilde{y}(s))\|_{\gamma(H,L^{2})}^{2}\right)\mathrm{d}s\bigg{]}.$
Using the Lipschitz condition Eq. 4.4 of $F$, we deduce
$\mathbb{E}\|e_{m}(t)\|_{L^{2}}^{2}\leqslant c\mathbb{E}\int_{0}^{t}\|e_{m}(s)\|_{L^{2}}^{2}\,\mathrm{d}s,\quad\forall t\in[0,T].$
Gronwall’s inequality then yields
$\sup_{t\in[0,T]}\mathbb{E}\|e_{m}(t)\|_{L^{2}}^{2}=0,\quad\forall m\geqslant 1.$
Since $e_{m}\in L_{\mathbb{F}}^{0}(\Omega;C([0,T];L^{4}))$, this implies $\|e_{m}\|_{C([0,T];L^{4})}=0$ holds $\mathbb{P}$-almost surely. Consequently,
$y=\widetilde{y}\quad\text{in }C([0,\sigma_{m}];L^{4})\text{ holds }\mathbb{P}\text{-almost surely, }\forall m\geqslant 1.$
As $y,\widetilde{y}\in L_{\mathbb{F}}^{0}(\Omega;C([0,T];L^{4}))$, we have $\lim_{m\to\infty}\mathbb{P}(\sigma_{m}=T)=1$. Thus, $y=\widetilde{y}$ in $C([0,T];L^{4})$ holds $\mathbb{P}$-almost surely, proving uniqueness for the mild solution to Eq. 1.1.
@@
Step 3. Fix any $p,q\in(2,\infty)$. By the regularity assertion in Eq. 3.1 and the growth estimate Eq. 4.1 for $F$, we deduce that $y-y^{3}$ belongs to $L_{\mathbb{F}}^{p}(\Omega\times(0,T);L^{q})$ and $F(y)$ is an element of $L_{\mathbb{F}}^{p}(\Omega\times(0,T);\gamma(H,L^{q}))$. Hence, the application of Lemma 4.3, in conjunction with Lemma 4.1, gives that
@@
while Lemma 4.4 implies $S_{1}F(y)\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{1,q})$. Additionally, the initial condition $y_{0}\in L^{\infty}$ ensures by Lemma 4.2 that $G\in L^{p}(0,T;\dot{H}^{2/p-\epsilon,q})$ for all $\epsilon>0$. Consequently, by the mild solution representation Eq. 2.7, we arrive at
@@
Finally, combining Eqs. 2.7, 4.17 and 4.18 establishes the regularity assertions in Eqs. 3.4 and 3.5. Moreover, the regularity assertion in Eq. 3.2 follows directly from Eq. 3.4 and the observation that $G\in C([0,T];L^{q})$, which is justified by the initial condition $y_{0}\in L^{\infty}$ and Lemma 4.2. This completes the proof of Theorem 3.1.
@@
We begin by summarizing several well-known results. It is known from [9] that
$\|P_{h}\|_{\mathcal{L}(L^{q})}\text{ is uniformly bounded in }h\text{ for all }q\in[1,\infty].$ (4.19)
This uniform stability of $P_{h}$ implies the following approximation property:
@@
Furthermore, it is a standard result (see, e.g., [13, Equations (12) and (13)]) that
$\lVert{I-\Delta_{h}^{-1}P_{h}\Delta}\rVert_{\mathcal{L}(\dot{H}^{2,q},L^{q})}$ $\leqslant ch^{2},\quad\forall\,q\in(1,\infty),$ (4.21)
$\lVert{I-\Delta_{h}^{-1}P_{h}\Delta}\rVert_{\mathcal{L}(\dot{H}^{1,q},L^{q})}$ $\leqslant ch,\quad\forall\,q\in[2,\infty).$ (4.22)
By applying complex interpolation between (4.21) and (4.22), one obtains
$\lVert{I-\Delta_{h}^{-1}P_{h}\Delta}\rVert_{\mathcal{L}(\dot{H}^{\alpha,q},L^{q})}\leqslant ch^{\alpha},\quad\forall\,\alpha\in[1,2),\quad\forall\,q\in[2,\infty).$ (4.23)
Combining this with Eq. 4.20 via the triangle inequality yields
$\lVert{P_{h}-\Delta_{h}^{-1}P_{h}\Delta}\rVert_{\mathcal{L}(\dot{H}^{\alpha,q},L^{q})}\leqslant ch^{\alpha},\quad\forall\alpha\in[1,2],\,\forall q\in[2,\infty),$ (4.24)
which in turn implies the estimate
$\lVert{P_{h}\Delta^{-1}-\Delta_{h}^{-1}P_{h}}\rVert_{\mathcal{L}(\dot{H}^{\alpha-2,q},L^{q})}\leqslant ch^{\alpha},\quad\forall\alpha\in[1,2],\,\forall q\in[2,\infty).$ (4.25)
Remark 4.2.
The projection operator $P_{h}$ can be extended to act as a linear operator from $\dot{H}^{-1,q}$ to $X_{h}$ for any $q\in(1,\infty)$. This extension is defined via the identity
$\int_{\mathcal{O}}(P_{h}v)v_{h}\,\mathrm{d}x=\langle{v,v_{h}}\rangle,$
[…]
Next, we introduce the discrete Sobolev space and recall some standard results. Let $q\in(1,\infty)$. For any $\alpha\in\mathbb{R}$, define $\dot{H}_{h}^{\alpha,q}$ as the space $X_{h}$ equipped with the norm
$\|v_{h}\|_{\dot{H}_{h}^{\alpha,q}}:=\|(-\Delta_{h})^{\alpha/2}v_{h}\|_{L^{q}},\quad v_{h}\in X_{h}.$
According to Theorem 3.1 (ii) in [19], the purely imaginary powers of $-\Delta_{h}$ are bounded in $\mathcal{L}(\dot{H}_{h}^{0,q})$ uniformly in $h$. As a consequence, an application of Theorem 4.17 from [24] yields, for any $\alpha\in(0,2)$, the equivalence
$c_{0}\|v_{h}\|_{\dot{H}_{h}^{\alpha,q}}\leqslant\|v_{h}\|_{[\dot{H}_{h}^{0,q},\dot{H}_{h}^{2,q}]_{\alpha/2}}\leqslant c_{1}\|v_{h}\|_{\dot{H}_{h}^{\alpha,q}},\quad v_{h}\in X_{h},$
with constants $c_{0},c_{1}>0$ independent of $h$. Here, $[\dot{H}_{h}^{0,q},\dot{H}_{h}^{2,q}]_{\alpha/2}$ denotes the complex interpolation space between $\dot{H}_{h}^{0,q}$ and $\dot{H}_{h}^{2,q}$; see [24, Chapter 2] for details. Given the standard inverse estimate
$\|v_{h}\|_{\dot{H}_{h}^{2,q}}\leqslant ch^{-2}\|v_{h}\|_{L^{q}},\quad v_{h}\in X_{h},$
complex interpolation yields the following general inverse estimate:
$\|v_{h}\|_{\dot{H}_{h}^{\beta,q}}\leqslant ch^{\alpha-\beta}\|v_{h}\|_{\dot{H}_{h}^{\alpha,q}},\quad v_{h}\in X_{h},\,\forall 0\leqslant\alpha<\beta<\infty.$ (4.26)
Applying this inverse estimate to Eq. 4.25, we obtain the estimate
$\|P_{h}\Delta^{-1}-\Delta_{h}^{-1}P_{h}\|_{\mathcal{L}(\dot{H}^{\alpha-2,q},\dot{H}_{h}^{\beta,q})}\leqslant ch^{\alpha-\beta}\quad\text{for all }\alpha\in[1,2],\,\beta\in[0,1],\,q\in[2,\infty).$ (4.27)
In analogy with Lemma 4.1, we also have the following $h$-uniform embedding result.
Lemma 4.7.
Let $0\leqslant\beta<\gamma\leqslant 1$ and $p,q\in(1,\infty)$. Then
$\|v_{h}\|_{\dot{H}_{h}^{\beta,q}}\leqslant c\|v_{h}\|_{(\dot{H}_{h}^{0,q},\dot{H}_{h}^{2,q})_{\gamma/2,p}},\quad v_{h}\in X_{h}.$
Additionally, we establish the following $h$-uniform Sobolev-type embeddings into $L^{\infty}$, whose proofs are given in Appendix A.
Lemma 4.8.
The following continuous embeddings hold:
  1. (i)
    If $q\in(3/2,\infty)$, then $\|v_{h}\|_{L^{\infty}}\leqslant c\|v_{h}\|_{\dot{H}_{h}^{2,q}}$ for all $v_{h}\in X_{h}$.
  2. (ii)
    If $\alpha\in(0,2)$ and $q\in(3/\alpha,\infty)\cap[2,\infty)$, then $\|v_{h}\|_{L^{\infty}}\leqslant c\|v_{h}\|_{\dot{H}_{h}^{\alpha,q}}$ for all $v_{h}\in X_{h}$.
Now, we introduce the discrete semigroup $e^{t\Delta_{h}}$ and the maximal $L^{p}$-regularity estimates associated with this semigroup. Let $e^{t\Delta_{h}}$, for $t\in[0,\infty)$, denote the analytic semigroup generated by the discrete Laplace operator $\Delta_{h}$. The following fundamental properties hold, with proofs provided in Appendix A.
Lemma 4.9.
The following estimates hold:
  • (i)
    For any $\alpha\in[0,1]$, $q\in(1,\infty)$, and $t>0$,
@@
  • (ii)
    For any $\alpha\in[0,2]$, $q\in(1,\infty)$, and $t>0$,
@@
For a function $g_{h}\in L^{1}(0,T;\dot{H}_{h}^{0,2})$, the discrete convolution operator $S_{0,h}$ is defined by
$(S_{0,h}g_{h})(t):=\int_{0}^{t}e^{(t-s)\Delta_{h}}g_{h}(s)\,\mathrm{d}s,\quad t\in[0,T].$ (4.30)
For a stochastic process $g_{h}\in L_{\mathbb{F}}^{2}(\Omega\times(0,T);\gamma(H,\dot{H}_{h}^{0,2}))$, the discrete stochastic convolution operator $S_{1,h}$ is given by
$(S_{1,h}g_{h})(t):=\int_{0}^{t}e^{(t-s)\Delta_{h}}g_{h}(s)\,\mathrm{d}W_{H}(s),\quad t\in[0,T].$ (4.31)
These operators satisfy the following fundamental discrete maximal $L^{p}$-regularity estimates.